# An Adjoint Formulation of Energetic Particle Confinement

### v1

@@ IV Escape Time
The phase space distribution of fast ion escape times can be conveniently illustrated by taking slices of the $\left(R,Z,\xi\right)$ of the PINN’s solution. The escape time of ions born with $\theta=0$ is shown in Fig. 6. Here, ions in the edge of the plasma can be lost quickly due to direct orbit loss, with ions initially located deeper in the plasma having much longer confinement times. Further noting that the white contour indicates the trapped-passing boundary, it is evident that co-current passing ions (positive pitches above the white contour) have far better confinement in comparison to counter-current ions. Similarly, in the trapped region, ions that begin with $\xi>0$ are much better confined compared to ions with $\xi<0$. Comparing the left and right panels of Fig. 2, this difference in confinement results from trapped ions born with $\xi<0$ at $\theta=0$ having orbits that extend to larger radii [Fig. 2(b)], in contrast to trapped ions born with $\xi>0$ [Fig. 2(a)]. The ions with the worst confinement are those just inside the trapped-passing boundary and $\xi<0$, which are expected to have the largest banana width, and thus be most susceptible to orbit loss.
@@
Figure 7 shows time slices of major radius $R$ and vertical height $Z$ for different pitches. These slices allow for the difference in confinement of co and counter current ions to be investigated more clearly. Considering passing co-current ions ($\xi=0.8$), these ions will propagate in the counter-clockwise direction in Fig. 7(a). Noting that ion orbits drift downward, this will result in ions that are initially on the outboard side not being on direct loss orbits, in contrast to those born on the inboard side. This leads to a notable in-out asymmetry of the fast ion population, as evident in Fig. 7(a). This trend is reversed for passing ions that begin in the counter-current direction [Fig. 7(d)]. A more subtle feature is that counter-current ions that are not on direct loss orbits are more likely to be lost in comparison to co-current ions. This is evident by the orange region in Fig. 7(d). The relatively poor confinement of these counter-current ions is due to the collisional trapping of these ions results in the orbit exploring larger minor radii [compare the red and green curves in Fig. 2(b)]. Hence, passing counter-current ions that are within a banana width of plasma edge are far more susceptible to being lost to the plasma compared to passing co-current ions.
@@
Cross-cuts of the pitch and poloidal angle distribution of loss times at different radii are shown in Fig. 8. Here, when approaching $r/a=1$ a discontinuity appears in the solution. This is evident in Fig. 8(a), where for $\theta\approx 0$ and $\theta\approx\pi$, the solution changes rapidly when traversing these discontinuous regions. The origin of this discontinuity is that for ions in the lower half of the tokamak (i.e. $\theta<0$) and $r/a=1$, they are lost immediately. In contrast, ions located just above $\theta=0$ or $\theta=\pi$ are often not on direct loss orbits, and thus can be confined for a long time. For example, co-current ions just above $\theta=0$, will drift inward, and thus have a finite escape time, whereas those just below $\theta=0$ are lost immediately to the boundary. An analogous discontinuity in the solution is present for $\theta=\pi$. To relax these discontinuities, we have removed all training and test points from these regions, such that the PINN is able approximate these discontinuities with a sharply varying solution, as evident in Fig. 8(a). Moving radially away from the $r/a=1$ boundary, panels (b)-(d) of Fig. 8 show the solution at radial locations between $r/a=0.95$ and $r/a=0.8$, indicating vastly different escape times of ions initially located at different pitch and poloidal angles. The trapped-passing boundary is also indicated in this figures, with the escape time of the solution often varying sharply across the trapped-passing region.
@@
The escape time can also be computed by the particle-based solve JONTA described in Sec. III.2. Here, marker particles are initialized randomly across the spatial domain at a given pitch. Once a particle’s minor radius exceeds $r_{wall}$, its escape time and location is saved. By pushing a large number of marker particles, the average escape time of the marker particle can be computed. Noting the conceptual simplicity of this approach, we will use it to validate the predictions of the PINN. Figure 9 shows four slices in the $\left(R,Z\right)$ plane at constant pitch computed from JONTA. Here, ten million marker particles are randomly distributed at a given pitch, with their loss times recorded after they escape from the domain. These escape times are then binned, allowing for the average escape time to be computed over the $\left(R,Z\right)$ plane for a given pitch. The particles were evolved until $t_{final}=2\times 10^{6}$, a time at which less than 0.001 percent of the original particles remained in the domain. Four cross cuts are shown in Fig. 9. These four cross cuts are for the same physics parameters and pitches as Fig. 7 above. Good agreement between the average escape time predicted by JONTA and the PINN are evident. In particular, both approaches are able to recover the rapid escape of ions that are initiated on direct loss orbits (blue regions), ions lost due collision trapping/detrapping directly onto a loss orbit (yellow and green regions), and ions initially located near the magnetic origin that must slowly diffuse out (red regions). The JONTA simulations also reveal a limitation of our implementation of a PINN. In particular, while the PINN is able to distinguish the complex edge structure of confined versus unconfined orbits, with excellent agreement with predictions of JONTA, it struggles to accurately capture the very long escape time for ions initially located deep inside the plasma. Specifically, the dark red region evident in Fig. 9 of the JONTA data, is not present in the PINN’s solution (see Fig. 7), indicating roughly a factor of two discrepancy in predicted confinement for these 20 keV well confined ions. The reason for this quantitative inaccuracy is that the PINN struggles to simultaneously resolve the fast direct orbit dynamics together with the slow collisional transport of the core ions, though as indicated in Fig. 4, this discrepancy slowly shrinks as the model is trained. The resolution of this disparate physics will require either the inclusion of data in training the PINN, using data taken directly from JONTA for example, or modifications to the training of the PINN. These extensions will be discussed further below.
@@
Now considering an energetic ion population with $50\;\text{keV}$, similar physical trends are evident, but the larger ion orbits, lead to a larger number of ions beginning on direct loss orbits. In addition, the higher energy of the particles leads to a lower collision time, and thus energetic ions initially located in the core have a longer escape time. This latter property results in the problem exhibiting greater stiffness, with a faster bounce time but slower collision time, resulting in the PINN struggling to accurately resolve the escape time of the best confined orbits. This is evident in Fig. 10, where cross sections in the $\left(R,Z\right)$ plane computed both with JONTA and the PINN are compared. While the PINN fails to capture the precise escape time of the best confined particles, it does succeed in accurately capturing the phase space structure of the escape time. It thus acts as an effective tool for delineating well confined regions of phase space, from those where ions are directly lost, or lost after a small number of collisional scattering events. We note that while for the case of $20\;\text{keV}$ ions roughly a factor of two difference in the escape time was present for the best confined ions, for $50\;\text{keV}$ nearly an order of magnitude difference in the prediction of the best confined orbits is evident. This larger deviation in the prediction of the best confined ions is due to the increased time scale separation between the collisionless ion orbits and the slow collision time.
@@ V Conclusions
The inhomogeneous adjoint of the drift kinetic equation was used to evaluate the escape time of energetic ion populations. This quantity, computed in the limit of an axisymmetric tokamak plasma, captures fast ion losses due to direct orbit loss and collisional transport, thus providing a metric for energetic particle confinement. A PINN was used to evaluate the phase space dependence of the escape time, a challenging task due to the strong time scale separation between the fast bounce time of energetic ions, and their slow collision rate. While the offline training time of the PINN is substantial, the online inference time is rapid, typically microseconds per prediction. Although the PINN failed to quantitatively capture the escape time of the best confined ion orbits, it is able to robustly recover the phase space structure of the escape time, and thus acts as a means of distinguishing well confined orbits versus those on direct loss orbits, or those that are directly lost after being scattered into or out of the trapped or passing regions. We also note that the present analysis did not include fast ion slowing down, where the slowing down time for the 20 and 50 keV ions considered in this work would be far shorter than the escape times of the best confined orbits.
@@
For this proof-of-principle demonstration of the proposed framework we considered several simplifications to the complete system. Perhaps the primary simplifications were the assumption of (i) an axisymmetric magnetic geometry, (ii) a simplified collision operator, and (iii) idealized plasma profiles and shaping. Regarding item (i), since PINNs to date have not been used to solve the drift kinetic equation, beginning with an axisymmetry geometry was a logical starting point for assessing this approach’s strengths and weaknesses. We anticipate extending the present approach in future work to 3D magnetic geometry, though such an extension will likely require incorporating particle data into the training of the neural network. The particle-based solver JONTA will provide a convenient tool through which such data can be generated, where by providing an accurate calculation of the escape time at specific phase space locations, these may be included in the first term of the loss function defined by Eq. (10). By providing high fidelity data points at internal locations, for example in regions with large residuals or for the best confined ion orbits where the present implementation of a PINN has struggled, this will likely expedite the training of the PINN while providing a systematic means of improving its accuracy. Our aim will be to leverage data as a means of both treating more challenging scenarios, such as non-axisymmetric geometry, along with facilitating the learning of a parametric solution to the drift kinetic equation, enabling the escape time to be evaluated for a broad range of plasma conditions.
@@
An additional simplification was the use of a Lorentz collision operator, which does not account for the slowing down of energetic particles, a quantity of critical interest when assessing the amount of energy deposited by energetic ions into the bulk plasma. Despite this limitation, including Lorentz collisions allows for the collisional transport of fast ions to be assessed. Thus, while not providing an accurate indication of the escape time for well confined energetic particles, since such particles will likely slow down substantially during their long confinement time, it does provide a means of rapidly assessing particles on direct loss orbits, or those that are directly lost after being collisionally scattered into or out of the trapped particle region. The predictions of the present work will thus be most relevant to the outer region of the tokamak, where energetic particles are lost on relatively short time scales. Finally, homogeneous plasma profiles and a circularly shaped plasma were assumed in the present work. This is done for convenience, where we don’t anticipate including more realistic plasma profiles to pose a substantial challenge. Furthermore, since PINNs are mesh free, and thus well suited to treating shaped magnetic geometry van Milligen et al. (1995); Kaltsas and Throumoulopoulos (2022); Jang et al. (2024), we do not expect the treatment of shaped axisymmetric geometry to pose a major hurdle.
@@
Due to the strong time scale separation, evaluating fast ion transport is computationally intensive, creating a bottleneck in many optimization workflows. For the JONTA simulations presented in this paper, the shortest simulation required one and a half hours on a single NVIDIA Blackwell 200 GPU, with the longest requiring slightly over a day. Noting that each JONTA simulation was for a fixed value of pitch, computing four distinct pitches (Figs. 9 and 10) thus required several hours to days of computing time. A PINN, which encompasses all values of pitch, and thus only a single PINN must be trained, required approximately two days to train, where approximate results obtained after the SOAP phase of training typically take a few hours. Both approaches are thus computationally intensive, where while the PINN is able to recover smooth predictions of the escape time, it is unable to accurately compute the best confined ions. JONTA, in contrast, requires roughly ten million markers to obtain a relatively smooth, well resolved solution. We note that while the PINN also used a single NVIDIA Blackwell 200 GPU, the implementation of the SSBroyden optimizer does not make efficient use of the GPU, and thus was mostly trained using a CPU, which was the primary bottleneck in the training of the PINN.
@@
We also note that while the present paper has focused on the escape time of energetic ions several straightforward extensions are possible. One such extension would be to treat a broader range of fast ion metrics. Metrics of interest include the probability that an ion impacts a specific region of the wall from a given initial phase space location Särkimäki (2020), or the slowing down probability, which corresponds to the probability that an ion slows down to the thermal energy. Both of these fast ion metrics can be straightforwardly constructed from the steady state homogeneous adjoint equation by applying appropriate boundary conditions. The evaluation of this broader class of fast ion metrics will be left for future work. Finally, while we have considered energetic ions in the present formulation, extensions to the relativistic drift kinetic equation describing runaway electron (RE) evolution and transport can be straightforwardly developed. Such extensions will allow for ‘neoclassical’ corrections to RE generation rates to be inferred Rosenbluth and Putvinski (1997); McDevitt and Tang (2019); Arnaud and McDevitt (2024), or account for spatial transport due to processes such as the Ware pinch Nilsson et al. (2015); McDevitt et al. (2019). While we anticipate that the extremely short bounce time of REs compared to their long collision times will pose a substantial challenge, we note that the impact of ‘neoclassical’ effects on RE generation rates are most important near the critical energy for electrons to run away McDevitt and Tang (2019); Arnaud and McDevitt (2024), an energy that often ranges between several keV to tens of keV during a tokamak disruption. At these modest energies, the time scale separation between the transit or bounce time of a RE and the collision time will be drastically reduced. The extension of the present framework to REs will be pursued in future work.

### v2

@@ IV.2 Phase Space Structure of the Mean Escape Time
The phase space distribution of the fast ion mean escape time can be conveniently illustrated by taking slices in the three-dimensional $\left(R,Z,\xi\right)$ space. The mean escape time of ions born with $\theta=0$ (weak field side) is shown in Fig. 7 on a $\log_{10}$ scale. Here, counter-current ions ($\xi<0$) born in the edge of the plasma are lost quickly due to direct orbit loss, with both co and counter current ions initially located deeper in the plasma having much longer mean escape times. Further noting that the white contour indicates the trapped-passing boundary, it is evident that co-current passing ions (positive pitches above the white contour) have far better confinement in comparison to counter-current ions. Similarly, in the trapped region, ions that begin with $\xi>0$ are much better confined compared to ions with $\xi<0$. Comparing the left and right panels of Fig. 3, this difference in confinement results from trapped ions born with $\xi<0$ at $\theta=0$ having orbits that extend out to larger radii [Fig. 3(b)], in contrast to trapped ions born with $\xi>0$ [Fig. 3(a)]. The ions with the worst confinement are those just inside the trapped-passing boundary and $\xi<0$, which are expected to have the largest banana width, and thus be most susceptible to orbit loss.
@@
Figure 8 shows slices of major radius $R$ and vertical height $Z$ for different pitches. These slices allow for the difference in confinement of co and counter current ions to be investigated more clearly. Considering passing co-current ions ($\xi=0.8$), these ions will propagate in the counter-clockwise direction in Fig. 8(a). Noting that ion orbits drift downward, this will result in ions that are initially on the outboard side not being on direct loss orbits, in contrast to those born on the inboard side. This leads to a notable in-out asymmetry of the fast ion population, as evident in Fig. 8(a). This trend is reversed for passing ions that begin in the counter-current direction [Fig. 8(d)]. A more subtle feature is that counter-current ions that are not on direct loss orbits are more likely to be lost in comparison to co-current ions. This is evident by the orange region in Fig. 8(d). The relatively poor confinement of these counter-current ions is due to the collisional trapping of these ions resulting in the orbit exploring larger minor radii [compare the red and green curves in Fig. 3(b)]. Hence, passing counter-current ions that are within a banana width of the plasma edge are far more susceptible to being lost to the plasma compared to passing co-current ions.
@@
Cross-cuts of the pitch and poloidal angle distribution of loss times at different radii are shown in Fig. 9. Here, when approaching $r/a=1$ a discontinuity appears in the solution. This is evident in Fig. 9(a), where for $\theta\approx 0$ and $\theta\approx\pm\pi$, the solution changes rapidly when traversing these discontinuous regions. The origin of this discontinuity is that for ions in the lower half of the tokamak (i.e. $\pi<\theta<0$) and $r/a=1$, they are lost immediately. In contrast, ions located just above $\theta=0$ or $\theta=\pm\pi$ are often not on direct loss orbits, and thus can be confined for a long time. For example, co-current ions just above $\theta=0$, will drift inward, and thus have a finite escape time, whereas those just below $\theta=0$ are lost immediately to the boundary. An analogous discontinuity in the solution is present for $\theta=\pm\pi$ for counter-current ions. To relax these discontinuities, we have removed all training and test points from these regions, such that the PINN is able to approximate these discontinuities with sharply varying, but continuous solutions, as evident in Fig. 9(a). Moving radially away from the $r/a=1$ boundary, panels (b)-(d) of Fig. 9 show the solution at radial locations between $r/a=0.95$ and $r/a=0.8$, indicating vastly different mean escape times of ions initially located at different pitch and poloidal angles. In particular, the phase space region of prompt ion loss, which initially occupies roughly half of the $\left(\xi,\theta\right)$ domain for $r/a\approx 1$, shrinks to a narrow sliver just inside the trapped-passing boundary for $r/a=0.8$. This is indicated in Fig. 9(d), where the solid white curve indicates the trapped-passing region.
@@
IV.3 Comparison with Monte Carlo Solutions
The mean escape time can also be computed by the particle-based solver JONTA described in Sec. III.2. Here, marker particles are initialized randomly across the spatial domain at a given pitch. Once a particle’s minor radius exceeds $r_{wall}$, its escape time and location are saved. By pushing a large number of marker particles, the mean escape time can be computed as a function of the ion’s initial $R$ and $Z$. Noting the conceptual simplicity of this approach, we will use it to validate the predictions of the PINN. Figure 10 shows four slices in the $\left(R,Z\right)$ plane at constant pitch computed from JONTA. Here, ten million marker particles are randomly distributed at a given pitch, with their escape times recorded after they escape from the domain. These escape times are then binned, allowing for the mean escape time to be computed over the $\left(R,Z\right)$ plane for a given pitch. The particles were evolved until $t_{final}=2\times 10^{6}$, a time at which nearly all of the original particles remained in the domain. Four cross cuts are shown in Fig. 10. These four cross cuts are for the same physics parameters and pitches as Fig. 8 above. Good agreement between the mean escape time predicted by JONTA and the PINN are evident. In particular, both approaches are able to recover the rapid escape of ions that are initiated on direct loss orbits (blue regions), ions lost due to collisional trapping/detrapping directly onto a loss orbit (yellow and green regions), and ions initially located near the magnetic origin that must slowly diffuse out (red regions). The JONTA simulations also reveal a limitation of our implementation of a PINN. In particular, while the PINN is able to distinguish the complex edge structure of confined versus unconfined orbits, with excellent agreement with predictions of JONTA, it struggles to accurately capture the very long mean escape time for ions initially located deep inside the plasma. Specifically, the dark red region evident in Fig. 10 of the JONTA data, is not present in the PINN’s solution (see Fig. 8), indicating roughly a factor of two and a half discrepancy in predicted confinement for these 20 keV well confined ions. The reason for this quantitative inaccuracy is that the PINN struggles to simultaneously resolve the fast direct orbit dynamics together with the slow collisional transport of the core ions, though as indicated in Fig. 5, this discrepancy slowly shrinks as the model is trained. The resolution of this disparate physics will require (i) the inclusion of data in training the PINN, e.g. using data taken directly from JONTA for example, (ii) allowing the PINN to train for more epochs which has been verified to slowly reduce the discrepancy, or (iii) modifications to the training of the PINN. These extensions will be discussed further below.
@@ IV.3 Comparison with Monte Carlo Solutions
Now considering an energetic ion population with $50\;\text{keV}$, similar physical trends are evident, but the larger ion orbits lead to a larger number of ions beginning on direct loss orbits (see Fig. 11). In addition, the higher energy of the particles leads to a lower collision time, and thus energetic ions initially located in the core have a longer escape time. This latter property results in the problem exhibiting greater stiffness, with a faster bounce time but slower collision time, resulting in the PINN struggling to accurately resolve the escape time of the best confined orbits. This is evident in Fig. 11, where cross sections in the $\left(R,Z\right)$ plane computed both with JONTA and the PINN are compared. While the PINN fails to capture the precise mean escape time of the best confined particles, it does succeed in accurately capturing the phase space structure of the mean escape time across the edge region, where ions have the shortest mean escape times. It thus acts as an effective tool for delineating well confined regions of phase space, from those where ions are directly lost, or lost after a small number of collisional scattering events. We note that while for the case of $20\;\text{keV}$ ions a factor of roughly two and a half difference in the mean escape time was present for the best confined ions, for $50\;\text{keV}$ roughly a factor of five difference in the prediction of the best confined orbits is evident. This larger deviation in the prediction of the best confined ions is due to the increased time scale separation between the collisionless ion orbits and the slow collision time. As discussed further in Sec. V, while further training or hyperparameter tuning can likely reduce these discrepancies, a robust means of improving the predictions of the PINN will be the incorporation of small quantities of data.
@@ V Conclusions
The inhomogeneous adjoint of the drift kinetic equation was used to evaluate the mean escape time of energetic ion populations. This quantity, computed in the limit of an axisymmetric tokamak plasma, captures fast ion losses due to direct orbit loss and collisional transport, thus providing a metric for energetic particle confinement. A PINN was used to evaluate the phase space dependence of the mean escape time, a challenging task due to the strong time scale separation between the fast bounce time of energetic ions, and their slow collision rate. While the offline training time of the PINN is substantial, the online inference time is rapid, typically microseconds per prediction. Although the PINN failed to quantitatively capture the mean escape time of the best confined ion orbits, it is able to robustly recover the phase space structure of the mean escape time, and thus acts as a means of distinguishing well confined orbits and phase space regions of prompt ion loss. We also note that the present analysis did not include fast ion slowing down, where the slowing down time for the 20 and 50 keV ions considered in this work would be far shorter than the mean escape times computed in this work for the best confined orbits.
@@
For this proof-of-principle demonstration of the proposed framework we considered several simplifications to the complete system. Perhaps the primary simplifications were the assumption of (i) an axisymmetric magnetic geometry, (ii) a simplified collision operator, and (iii) idealized plasma profiles and shaping. Regarding item (i), since PINNs to date have not been used to solve the drift kinetic equation, beginning with an axisymmetry geometry was a logical starting point for assessing this approach’s strengths and weaknesses. We anticipate extending the present approach in future work to 3D magnetic geometry Law et al. (2022), though such an extension will likely require incorporating particle data into the training of the neural network. The particle-based solver JONTA will provide a convenient tool through which such data can be generated, where by providing an accurate calculation of the mean escape time at specific phase space locations, these may be included in the first term of the loss function defined by Eq. (10). By providing high fidelity data points at internal locations, for example in regions with large residuals or for the best confined ion orbits where the present implementation of a PINN has struggled, this will likely expedite the training of the PINN while providing a systematic means of improving its accuracy. Our aim will be to leverage data as a means of both treating more challenging scenarios, such as non-axisymmetric geometry, along with facilitating the learning of a parametric solution to the drift kinetic equation, enabling the mean escape time to be evaluated for a broad range of plasma conditions.
@@
An additional simplification was the use of a Lorentz collision operator, which does not account for the slowing down of energetic particles, a quantity of critical interest when assessing the amount of energy deposited by energetic ions into the bulk plasma. Despite this limitation, including Lorentz collisions allows for the collisional transport of fast ions to be assessed. Thus, while not providing an accurate indication of the mean escape time for well confined energetic particles, since such particles will likely slow down substantially during their long confinement time, it does provide a means of rapidly assessing particles on direct loss orbits, or those that are directly lost after being collisionally scattered into or out of the trapped particle region. The predictions of the present work will thus be most relevant to the outer region of the tokamak, where energetic particles are lost on relatively short time scales. Finally, homogeneous plasma profiles and a circularly shaped plasma were assumed in the present work. This was done for convenience, where we anticipate including more realistic plasma profiles to not pose a substantial challenge. Furthermore, since PINNs are mesh free, and are thus well suited to treating shaped magnetic geometry van Milligen et al. (1995); Kaltsas and Throumoulopoulos (2022); Jang et al. (2024), we do not expect the treatment of shaped axisymmetric geometry to pose a major hurdle. Similarly, strong radial electric fields that emerge in the pedestal region Brzozowski et al. (2019) along with realistic equilibrium density and temperature profiles could easily be incorporated.
@@
Due to the strong time scale separation, evaluating fast ion transport is computationally intensive, creating a bottleneck in many optimization workflows. For the JONTA simulations presented in this paper, the shortest simulation required one and a half hours on a single NVIDIA Blackwell 200 GPU, with the longest requiring slightly over a day. Noting that each JONTA simulation was for a fixed value of pitch, computing four distinct pitches (Figs. 10 and 11) thus required several hours to days of computing time. A PINN, which encompasses all values of pitch, and thus only a single PINN must be trained, required approximately three days to train, where approximate results obtained after the SOAP phase of training typically take a few hours. Both approaches are thus computationally intensive, where while the PINN is able to recover smooth predictions of the mean escape time, it is unable to accurately compute the best confined ions. JONTA, in contrast, requires roughly ten million markers to obtain a relatively smooth, well resolved solution. We note that while the PINN also used a single NVIDIA Blackwell 200 GPU, the implementation of the SSBroyden optimizer does not make efficient use of the GPU, and thus was mostly trained using a CPU, which was the primary bottleneck in the training of the PINN.
@@
We also note that while the present paper has focused on the mean escape time of energetic ions, several straightforward extensions are possible. One such extension would be to treat a broader range of fast ion metrics. Metrics of interest include the probability that an ion impacts a specific region of the wall from a given initial phase space location Särkimäki (2020); Salewski et al. (2011), or the slowing down probability, which corresponds to the probability that an ion slows down to the thermal energy. Both of these fast ion metrics can be straightforwardly constructed from the steady state homogeneous adjoint equation by applying appropriate boundary conditions. The evaluation of this broader class of fast ion metrics will be left for future work. Finally, while we have considered energetic ions in the present formulation, extensions to the relativistic drift kinetic equation describing runaway electron (RE) evolution and transport can be developed following an analogous approach. Such extensions will allow for ‘neoclassical’ corrections to RE generation rates to be inferred Rosenbluth and Putvinski (1997); McDevitt and Tang (2019); Arnaud and McDevitt (2024), or account for spatial transport due to processes such as the Ware pinch Nilsson et al. (2015); McDevitt et al. (2019). While we anticipate that the extremely short bounce time of REs compared to their long collision times will pose a substantial challenge compared to the purely momentum space treatment employed in Refs. Arnaud et al. (2024); McDevitt et al. (2025), we note that the impact of ‘neoclassical’ effects on RE generation rates are most important near the critical energy for electrons to run away McDevitt and Tang (2019); Arnaud and McDevitt (2024), an energy that often ranges between several keV to tens of keV during a tokamak disruption. At these modest energies, the time scale separation between the transit or bounce time of a RE and the collision time will be drastically reduced. The extension of the present framework to REs will be pursued in future work.